the sum of the digits of a two digit number is 10. The number formed by reversing the digits is 4 less than 5 times the original number. Find the original number.
can you explain this one to me thank you so much.

Answer by lwsshak3(11628)

the sum of the digits of a two digit number is 10. The number formed by reversing the digits is 4 less than 5 times the original number. Find the original number.
can you explain this one to me thank you so much.
**
let u=units digit
let t=tens digit
u+t=10
t=10-u
original number=10t+u
reversed digit number=10u+t
..
10u+t=5(10t+u)-4
10u+t=50t+5u-4
49t-5u=4
49(10-u)-5u=4
490-49u-5u=4
54u=486
u=486/54=9
t=10-u=1
original number=10t+u=10+9=19
..
check:
reversed digit number=10u+t=90+1=91
5 times the original number=5*19=95
difference=4